### v2

# TRAWL: Tensor Reduced and Approximated Weights for Large Language Models

Abstract
Recent research has shown that pruning large-scale language models for inference is an effective approach to improving model efficiency, significantly reducing model weights with minimal impact on performance. Interestingly, pruning can sometimes even enhance accuracy by removing noise that accumulates during training, particularly through matrix decompositions. However, recent work has primarily focused on single matrix decompositions or lower precision techniques, which may fail to fully capture structural patterns. To address these limitations, we introduce TRAWL (Tensor Reduced and Approximated Weights for Large Language Models), a technique that applies tensor decomposition across multiple weight matrices to effectively denoise LLMs by capturing global structural patterns. Our experiments show that TRAWL improves model performance by up to 16% over baseline models on benchmark datasets, without requiring additional data, training, or fine-tuning.
TRAWL: Tensor Reduced and Approximated Weights
for Large Language Models
Yiran Luo1*, Het Patel1*, Yu Fu1, Dawon Ahn1, Jia Chen2, Yue Dong1, Evangelos E. Papalexakis1 1 Department of Computer Science and Engineering, UC Riverside, USA 1 {yluo147, hpate061, yfu093, dahn017, yued, epapalex}@ucr.edu 2 Department of Electrical and Computer Engineering, UC Riverside, USA 2 jiac@ucr.edu * These authors contributed equally to this work.
1 Introduction
The success of Transformer-based models like the GPT series Radford et al. (2018, 2019); Brown et al. (2020), and LLaMA Touvron et al. (2023) is often attributed to their large scale Kaplan et al. (2020). While the large number of parameters allows these models to capture complex patterns, training and deploying them requires significant computational resources and energy Samsi et al. (2023); Zhao et al. (2022), making them impractical for many real-world applications.
To address these limitations, recent studies have explored post-training compression techniques, showing minimal impact on inference performance Han et al. (2015); Sharma et al. (2023); Saha et al. (2024). Interestingly, some research suggests that certain pruning methods can reduce computation time while also enhancing performance. For instance, LASER Sharma et al. (2023) applies Singular Value Decomposition (SVD) on weight matrices for model compression, demonstrating that removing noise introduced during training can improve performance. CALDERA Saha et al. (2024) builds on LASER by further compressing weight matrices into lower precision formats, achieving significant size reductions and lower computational costs. However, both methods operate on individual matrices, potentially missing broader structural patterns across the model.
In this paper, we introduce TRAWL (Tensor Reduced and Approximated Weights for LLMs), a novel approach that extends denoising by leveraging tensor decomposition across multiple weight matrices. Unlike methods that focus on individual matrices, TRAWL stacks multiple weight matrices into higher-order tensors, enabling more effective exploitation of the model’s inherent structures.
Our experiments on two Transformer-based models- RoBERTa Liu et al. (2019) and GPT-J Wang and Komatsuzaki (2021) across three datasets (BigBench WikiQA Srivastava et al. (2022), BiosProfession De-Arteaga et al. (2019), and HotpotQA Yang et al. (2018)) demonstrate that TRAWL consistently outperforms single matrix decomposition approaches, with accuracy improvements of up to 16%. By leveraging advanced tensor decomposition, TRAWL effectively reduces noise introduced during training and enhances performance without requiring additional training or fine-tuning. These results highlight TRAWL as a practical and efficient optimization technique for LLMs, delivering significant accuracy gains over baseline models.
Our key contributions are:
We propose a post-training compression approach that uses tensor decomposition across multiple weight matrices, enabling more effective noise reduction and rank compression of model weights.
We conducted extensive experiments on two models across three datasets, showing that TRAWL improves accuracy and generalization over existing methods without additional training or fine-tuning, making it highly practical for real-world applications. All code from this study will be publicly available upon publication to support transparency and future research.111https://anonymous.4open.science/r/TRAWL
2 Related Work
The two most relevant works to ours are LAyer SElective Rank Reduction (LASER) Sharma et al. (2023) and CALDERA Saha et al. (2024), which leverage matrix decomposition and low-rank approximation to compress LLMs. LASER applies singular value decomposition (SVD) to reduce weight matrix rank, improving performance by removing noise. CALDERA combines low-precision and low-rank matrices to reduce memory usage while maintaining accuracy.
Other tensor-based methods, such as Low-Rank Economic Tensor-Train Adaptation (LoRETTA) Yang et al. (2024) significantly reduce trainable parameters through tensor-train decomposition, while techniques like Tensor Contraction Layers (TCLs) Kossaifi et al. (2017) incorporate tensor contractions as end-to-end trainable neural network layers to reduce the dimensionality of the activation tensors offering a way to reduce model parameters.
Building on these approaches, TRAWL fully leverages tensor structures in LLMs by applying tensor decomposition across multiple weight matrices, achieving performance improvements without additional training or fine-tuning.
3 Methodology
This section provides background on tensor decomposition and describes our proposed method.
3.1 Tensor Decomposition
Tensor decomposition is a technique for analyzing multi-dimensional data by breaking a tensor into lower-dimensional components Sidiropoulos et al. (2017). A tensor, essentially a multi-dimensional array, can be decomposed using various tensor decomposition methods. In this paper we focus on two: CANDECOMP/PARAFAC (CP) decomposition and Tucker decomposition.
CP decomposition represents a tensor as a sum of rank-one tensors. For a tensor of order three, this is expressed as:
where , , and are factor vectors, denotes the outer product, and is the rank. is the low-rank approximation of and it represents the original tensor as a sum of a few rank-one tensors. This reduces storage and computation while preserving essential information.
In the case of Tucker decomposition, is approximated as a core tensor multiplied by factor matrices , , and along each mode:
where denotes the -mode product and essentially multiplies all the slices of the -th mode of the tensor with the corresponding factor matrix.
3.2 Proposed Approach
| Model | Approach | BigBench WikiQA | BiosProfession | HotpotQA | |||
| Accuracy | Loss | Accuracy | Loss | Accuracy | Loss | ||
| RoBERTa | Baseline | 28.0% | 9.07 | 64.5% | 4.91 | 6.1% | 10.99 |
| LASER | 30.7% | 7.69 | 72.5% | 6.44 | 6.7% | 10.53 | |
| 43.46% | 9.25 | 73.07% | 5.20 | 6.28% | 12.67 | ||
| 41.01% | 8.97 | 71.47% | 7.55 | 6.58% | 12.27 | ||
| GPT-J | Baseline | 51.84% | 3.52 | 75.58% | 4.64 | 19.6% | 3.40 |
| LASER | 65.9% | 2.86 | 82.1% | 4.91 | 19.5% | 3.39 | |
| 68.1% | 2.78 | 82.35% | 4.65 | 20.0% | 3.47 | ||
| 66.3% | 2.94 | 79.58% | 4.43 | 19.7% | 3.44 | ||
TRAWL focuses on tensor compression across multiple weight matrices, distinguishing itself from existing methods Sharma et al. (2023); Saha et al. (2024) by emphasizing how matrices are stacked for decomposition.
We propose an effective approach by stacking and decomposing weight matrices one layer at a time. In more details, we stack the selected matrices for each layer individually and apply tensor decomposition layer by layer. Specifically, for each layer , we form a 3-mode tensor :
Tensor decomposition is applied to each separately, and the impact on performance is evaluated layer by layer. The matrices in each layer are replaced by their decomposed counterparts and reverted after evaluation. This process is repeated for all layers. The approach is depicted in Figure 1.
Additionally, we explored two conventional tensor formulation methods (Appendix 7.2): stacking selected matrices across all layers and stacking by dividing the matrices into three-layer segments. Our results demonstrate that both approaches underperform compared to our one-layer-at-a-time strategy. A detailed experimental comparison is provided in Appendix 7.3.
4 Experiments and Results
4.1 Experimental Setups
We primarily focus on two models, GPTJ-6B Wang and Komatsuzaki (2021) and RoBERTa Liu et al. (2019), and tested them with three datasets, BigBench WikiQA Srivastava et al. (2022), BiosProfession De-Arteaga et al. (2019), and HotPotQA Yang et al. (2018). In our experiments, we applied two tensor decomposition methods—CP decomposition and Tucker decomposition—to better understand the effect of tensor decomposition on LLMs. These were compared with baseline experiments where no tensor decomposition was applied.
To apply tensor decomposition, we transpose the Fully Connected (FC) input or output matrix to align their shapes for stacking and decomposition. After decomposition and low-rank approximation, the matrix is transposed back to its original shape before being substituted into the model. For the Query, Key, Value, and Output from attention heads (QKVO) matrices, stacking is straightforward since they share the same shape, allowing easy decomposition. More details are in Appendix 7.1.
4.2 Results
Table 1 compares the performance of the baseline, LASER, and TRAWL methods for both RoBERTa and GPT-J models across three datasets: BigBench WikiQA, BiosProfession, and HotpotQA. Additional results from other tensorization strategies, including segmented layers and across-the-model experiments, are provided in Appendix 7.3.
First, we observe that TRAWL, our tensor decomposition approach across multiple weight matrices, consistently outperforms the baseline compression method, LASER. For the BigBench WikiQA dataset, TRAWLCP delivers the best accuracy, with 43.46% for RoBERTa and 68.1% for GPT-J, outperforming LASER and the baseline. It also achieves the lowest loss for GPT-J, while LASER has the lowest for RoBERTa. In the BiosProfession dataset, TRAWLCP continues to lead with 73.07% accuracy for RoBERTa and 82.35% for GPT-J, while improving GPT-J’s loss. For HotpotQA, the improvements are smaller, with TRAWLCP slightly ahead of LASER for GPT-J, and similar results across methods for RoBERTa.
In summary, TRAWL consistently outperforms the baseline and LASER methods, particularly with CP decomposition in the FC layers. The most significant improvements are seen with the GPT-J model, especially on the BigBench WikiQA and BiosProfession datasets. The observed benefits, however, are dataset-dependent, with HotpotQA showing only marginal improvements.
4.3 Ablations
To evaluate the effect of stacking and decomposing different layers, we compared the performance of CP-decomposed fully connected (FC) weight matrices across various layers of RoBERTa on multiple datasets. The most notable improvements occurred in the FC weight matrices of the final layers, where tensor decomposition led to significant accuracy gains. In contrast, decomposing the QKVO matrices yielded minimal benefits and, in some cases, even reduced performance. As shown in Figure 2, the later-stage FC layers exhibited the largest accuracy improvements, while earlier layers, particularly the QKVO matrices, showed little to no benefit, and in some instances, decreased accuracy. We also observe similar trends on GPT-J, where the results are presented in Appendix Figure 4.
Additionally, there are more fluctuations in accuracies in RoBERTa (Figure 2) compared to GPT-J, suggesting that RoBERTa’s performance is more sensitive to the specific rank and decomposition applied. These findings show that while tensor decomposition is beneficial, its impact is highly dependent on the specific layer, the corresponding weight matrix, and the tensor decomposition rank.
Last but not least, the one-layer-at-a-time approach usually resulted in the highest performance gains near the final few layers when decomposing FC weights. However, no significant improvements were observed for QKVO matrices. These results align with the findings of Sharma et al. (2023), highlighting that FC layers are key targets for effective compression. In this work we demonstrate that there exists meaningful low-rank structure when treating the FC weights jointly as a tensor. In the future, we aim to explore whether more complex models or alternative methods can yield improvements for QKVO matrices.
5 Conclusion
In this study, we introduce TRAWL, a novel method for denoising LLMs using tensor decomposition. TRAWL extracts specific weight matrices, stacks them into a tensor, and applies decomposition to replace the original matrices with low-rank approximations. Our results show that layer-by-layer decomposition and selective layer stacking significantly improve performance, with the most notable gains seen in the fully connected layers of the terminal layers. These findings highlight the effectiveness of adaptive decomposition techniques in enhancing LLM efficiency, sometimes even outperforming the original models through denoising. Future work will explore alternative decomposition methods, refine rank selection, and expand TRAWL’s application to a broader range of models and tasks, driving more efficient and accessible AI systems.
6 Limitations
While our study focused on tensor decomposition techniques like CANDECOMP/PARAFAC (CP) and Tucker, other methods such as Tensor training or Block-Term Decomposition (BTD) may offer further performance gains. Additionally, joint compression of heterogeneous layers forming "irregular tensors" presents another promising direction to explore for enhancing model efficiency.
Estimating the optimal rank for tensor decomposition remains a challenge, as there is no universal solution like the Singular Value Decomposition for matrices. Although we explored different ranks using available heuristics, there is still no "silver bullet." Refining these rank selection methods is an important area for future research.
Our method demonstrates that compressing certain layers and weight types can significantly improve performance. However, we lack a systematic, non-brute-force approach to determine the best combination of layers and weights for compression. This is a challenge shared by prior work, such as LASER, and remains an open problem for future exploration.
Low-rank compression provides two key benefits: reducing model size through approximation and improving performance via denoising. While we focused on the latter in this study, future research should investigate the practical systems implications, such as memory and compute savings for inference on various devices.
Finally, although our findings were consistent across two language models and three benchmark datasets, the extent to which they generalize to other models and datasets requires further validation. Future work should explore the broader applicability of our approach to ensure robust performance across diverse models and tasks.